OpenStudy (anonymous):

The sine of the complimentary angle to 75°

OpenStudy (cherry17ann):

This is English please go to math but to answer your question.... complement = ( 90 - x ) = ( 90 - 75 ) = 15 sine( complement of 75 ) = sin(15)